# AgentStudio: A Toolkit for Building General Virtual Agents

### v2

@@ 1 Introduction
Building general virtual agents that can utilize every software to enhance human productivity is a long-standing research challenge in artificial intelligence (Shi et al., 2017). Such agents perceive computer states as observations (e.g., screenshots) and take actions through human-computer interfaces (e.g., GUI, keyboard, and mouse) or function calling (e.g., APIs) in response to natural language instructions. Fueled by advancements in large language models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023), there has been impressive progress in virtual agents (Zheng et al., 2023; Zhang et al., 2024; Tan et al., 2024). However, deploying these agents in real-world settings requires satisfying three key desiderata: (1) achieving general grounding by interacting with both GUIs and APIs and handling complex observations like videos; (2) possessing extensive digital world knowledge to complete tasks across diverse applications; and (3) self-improving and generalizing to unseen situations through trial and error in open-ended environments. For example, an agent exporting charts from a spreadsheet and adding them to a video needs to interact precisely with GUI elements in both the spreadsheet and the video editor. Using API actions can speed up the process, and video observations can help monitor and make adjustments. The agent needs knowledge of these applications, e.g., understanding their workflows, to effectively perform the task. If the video editor’s layout is out of distribution, the agent should learn how to use it by interacting with the environment.
@@ 6 Related Work
Table 4: AgentStudio supports online interactions with real-world operating systems using the most generic observation (text (T), image (I), and video (V)) and action spaces (keyboard (K), mouse (M), and code) among the listed environments. It offers benefits such as user-friendly GUI/video annotation and task customization, lightweight installation, lightweight installation, and fine-grained datasets that decompose fundamental agent abilities.

### v3

@@ 1 Introduction
Building general virtual agents that can utilize every software to enhance human productivity is a long-standing research challenge in artificial intelligence (Shi et al., 2017). Such agents perceive computer states as observations (e.g., screenshots) and take actions through human-computer interfaces (e.g., GUI, keyboard, and mouse) or function calling (e.g., APIs) in response to natural language instructions. Fueled by advancements in large language models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023), there has been impressive progress in virtual agents (Zheng et al., 2023; Zhang et al., 2024; Tan et al., 2024). However, deploying these agents in real-world settings requires satisfying three key desiderata: (1) achieving general grounding by interacting with both GUIs and APIs and handling complex observations like videos, (2) possessing extensive digital world knowledge to complete tasks across diverse applications, and (3) self-improving and generalizing to unseen situations through trial and error in open-ended environments. For example, an agent exporting charts from a spreadsheet and adding them to a video needs to interact precisely with GUI elements in both the spreadsheet and the video editor. Using API actions can speed up the process, and video observations can help monitor and make adjustments. The agent needs knowledge of these applications, e.g., understanding their workflows, to effectively perform the task. If the video editor’s layout is out of distribution, the agent should learn how to use it by interacting with the environment.
@@ 6 Related Work
Table 4: AgentStudio supports online interactions with real-world operating systems using the most generic observation (text (T), image (I), and video (V)) and action spaces (keyboard (K), mouse (M), and code) among the listed environments. It offers benefits such as user-friendly GUI/video annotation and task customization, lightweight installation, and fine-grained datasets that decompose fundamental agent abilities.
@@ F.7 Additional Examples
F.8 Software Selection
Our software selection is based on systematic criteria. (1) Open-source: We select open-source software to avoid copyright issues and encourage community contributions. (2) Popularity: We include widely used tools, such as Google Workspace, VS Code, and LibreOffice, which are among the most prevalent applications in their respective categories. GIMP (GNU Image Manipulation Program) is one of the most popular open-source image editors. (3) Diversity: To represent a broad range of real-world use cases, we select software across different domains, including OS components, office suites, code editors, and image editors. Although our environment can support proprietary software like Photoshop, we prioritize using ones that do not require licenses for running benchmarks. Meanwhile, our benchmark tasks for GIMP are designed to be easily applicable to Photoshop as well, since most image-related tasks are evaluated by checking the output images.
F.9 More Technical Details
Our implementation uses Docker for a lightweight environment compared to virtual machines used by OSWorld (Xie et al., 2024). Key technical aspects include: (1) Performance & latency: The runtime mainly depends on the time for task reset, action generation & execution, auto-eval, and cleanup. AgentStudio is designed to mimic real-world computers operating in real time. The latency of reset, auto-eval, and cleanup depends on the Docker resource configuration, which can be improved by allocating additional resources to Docker. At present, the main bottleneck is the response time of LLM action generation, as well as the time required to execute LLM-generated code, rather than the latency in AgentStudio. (2) Scalability: Our architecture supports multiple instances running concurrently. The separation of the front end and back end ensures that LLM operations within Docker containers do not interfere with each other. Google API operations can also be isolated by setting up different accounts. (3) Reliability: Our Docker setup safeguards local files and systems from potential vulnerabilities. For sensitive operations, AgentStudio supports enabling human verification to enhance reliability and safety.